OpenStudy (anonymous):

The population of a town in 2000 was 430. The population is increasing at a rate of 0.9% every year. What will be the projected population of the town in 2010? Round your answer to the nearest whole number.

OpenStudy (anonymous):

Year 1 Pop ............ ...430 Year 1 increase(.9%)...430 x 0.009 = 3.87 Year 2 Pop ...............:433.87 Year 2 increase(9%)....433.87 x 0.009 = 3.90483 Year 3 Pop................437.77483 but, Pop in Yr 1 = 430 + 0.9% or 430(1+0.9%) = 433.87 Pop in Yr 2 = 430(1+0.9%)^2 = 437.77483 Therefore: Pop in Yr 10 = 430(1+0.9%)^10 = 470.3055653 Rounded to nearest whole number = 470